OpenStudy (anonymous):

The straight conductor below carries current i. Indicate the direction of the B field inside the rectangle and find an expression for the total flux  passing through the rectangular area A = bc

OpenStudy (anonymous):

If i = 2.0 A, c = 3 cm and b = 2*a, find the magnitude of the flux \[\phi \] through the rectangle

OpenStudy (anonymous):

Note: You must use integration PS: Do not bother please, if you are not familiar with such type of question

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

im not confident in systems of electricity. I believe flux in general is a calculation of the amount of flow thru a surface. if given a vector field F that defines the direction and magnitude of the flow in the system, then the surface gradient provides a normal (unit normal might need to be used) perpendicular direcion to be able to find the Force, or Flow, thru the surface.

OpenStudy (amistre64):

but im not too sure how i would implement that with the information given here

OpenStudy (anonymous):

\[\int\limits_{}^{}dΦr=\int\limits_{a}^{a+b}(μ0Ι/2πr)(cdr)\]

OpenStudy (anonymous):

\[Φr=μ0Ιc/2π\int\limits_{a}^{a+b}dr/r\]

OpenStudy (anonymous):

\[Φr=(μ0Ιc/2π)\ln (a+b/a)\]

OpenStudy (anonymous):

put the numbers,get your flux